OpenStudy (anonymous):

how do I turn (4x-1) into (1-4x)

OpenStudy (stamp):

\[-(4x-1) =-(4x)-(-1)=-4x+1=1-4x\] So basically (4x-1) = -(1-4x) and (1-4x) = -(4x-1)

OpenStudy (anonymous):

So multiply by -1

OpenStudy (stamp):

Yes.

OpenStudy (stamp):

(4x-1) * -1 = -(4x-1) = -4x + 1 = 1 - 4x

OpenStudy (anonymous):

Do I have to multiply a numerator too

OpenStudy (anonymous):

I have 24x^4-8x^3+3/(4x-1)^2 And on wolfram it simplifies to 24x^4-8x^3+3/(1-4x)^2

OpenStudy (stamp):

Depends on the problem. But if you multiply a numerator by -1, you must distribute a -1 to the denominator as well. Example: \[\frac{(4x-1)}{1-4x}=\frac{-(1-4x)}{1-4x} = -1\]

OpenStudy (stamp):

\[ (4x-1)^2 = (4x-1)(4x-1)=-1(1-4x)*-1(1-4x)\] \[=(-1)^2(1-4x)^2=(1-4x)^2\]